# Distilling Spectral Graph for Object-Context Aware Open-Vocabulary Semantic Segmentation

### v1

@@ 1 Introduction
Open-vocabulary semantic segmentation (OVSS) [28, 50, 9, 49, 53] aims to predict pixel-level labels for arbitrary prompts defined by user input (e.g., the proper noun Space Needle in Fig. 1(a)), advancing beyond previous models scoped to predefined static classes [35, 18, 10]. To achieve this, fully supervised OVSS models [11, 51, 48] enhance generalization to seen classes using labeled training data. Yet, their reliance on seen classes during training may risk overfitting and limit scalability, as labeled data is required for retraining to adapt to new domains [38].
@@ 2 Related Works
Open-vocabulary semantic segmentation (OVSS) aims to segment images corresponding to arbitrary user-given prompts, using vision-language aligned representations (e.g., CLIP [34]). Meanwhile, training-free approaches [3, 4, 25, 53] have risen for their ability to adapt directly to new tasks without extra training. Among these, methods [43, 53, 25, 16] that do not rely on additional datasets have gained attention, as they require no extra adaptation effort, aligning better with the objective of OVSS. Earlier approaches investigated leveraging CLIP features for better localization [53, 26] or for grounding proposed masks [42, 20]. SCLIP [43] uses correlative self-attention to improve inter-patch correlations. NACLIP [16] applies a Gaussian kernel to attention maps for better spatial consistency, and ProxyCLIP [25] integrates patch-wise consistent features from DINO [8]. Although these methods demonstrated strong results, they often struggle to group different parts of the same object due to a lack of explicit dense label knowledge. To this end, we attribute the underlying challenges of training-free OVSS by enriching the object-level contextual understanding.
@@
Object-Level Semantic Segmentation. Semantic segmentation groups object components that have similar semantics into a single mask, traditionally achieved through models trained on extensive human-labeled, pixel-wise annotations [18, 10, 35]. However, despite recent advancements in unsupervised semantic segmentation [17, 37] (USS) that do not rely on dense labels, a core challenge persists: individual object components are not consistently grouped into a single entity. To address this issue, recent USS studies [21, 45, 36] have incorporated object-level representation learning to improve the model’s ability to merge single objects into unified entities. However, these efforts have not been considered in the training-free OVSS task, which faces similar challenges in an unsupervised context. Therefore, we focus on enriching object-level context in training-free OVSS to achieve a more effective grouping of components within each object, enabling clearer and more consistent segmentation across diverse visual domains.
@@ 3.2 Spectral Object-Level Context Distillation
Our model CASS aims to capture such object-level contextual interaction within the image features to associate different parts of the same object. However, the limited ability of CLIP to capture semantic relationships between patches often neglects the object-level context, as illustrated in the left column of Fig. 3 (${A}_{\text{CLIP}}$). To address this limitation, we leverage VFM (e.g., DINO [8]), enabling a deeper object-level contextual understanding. Among VFM features, we utilize VFM attention key component ${K}_{\text{VFM}}\in\mathbb{R}^{h\times N\times D_{h}}$. Specifically, we treat the VFM attention adjacency ${A}_{\text{VFM}}={K}_{\text{VFM}}{K}_{\text{VFM}}^{\top}\in\mathbb{R}^{h\times N\times N}$ as a graph, where $h$ represents the number of heads, and transfer crucial graph patterns from ${A}_{\text{VFM}}$ to ${A}_{\text{CLIP}}$. However, a key question remains: “How do we account for the multi-head attention mechanisms in both VFM and CLIP?" In other words, “how do we match $i$-th VFM attention head ${A}_{\text{VFM}}^{i}$ with $j$-th CLIP attention head ${A}_{\text{CLIP}}^{j}$?" A straightforward approach is to distill the attention heads sequentially (i.e., $i=j$). Nevertheless, recent studies [54, 27] indicate that different attention heads focus on separate parts of the image. Thus, matching the optimal heads between VFM and CLIP is essential for precise attention distillation.
@@ 3.3 Object Presence-Driven Object-Level Context
where $\gamma$ controls the balance between $\hat{\mathcal{S}}$ and object presence prior. This enables the final segmentation map to capture object-level details, enabling precise, object-aware segmentation with accurate class assignments.
@@ 4.2 Evaluation Results
We evaluate the proposed model against existing training-free OVSS methods quantitatively and qualitatively. For fair comparisons, we refrain from using the mask refinement (e.g., PAMR [1] or DenseCRF [22]) and reproduce the results from [25], with a window size of $224\times 224$ to match the baselines. We mainly report the mean Intersection over Union (mIoU) metric to evaluate our model. Additionally, we incorporate average pixel accuracy (pAcc) to provide more comprehensive evaluations.
@@ 4.3 Ablation Study
Distance Metric for Graph Matching. We evaluate different distance metrics $\mathcal{D}_{W}$ used in Sec. 3.2.1, including Wasserstein Distance, Cosine Distance, KL Divergence, and Euclidean Distance. As shown in Fig. 8, the Wasserstein Distance shows the best results as it provides an intuitive measure of the overall shape of sorted eigenvalue distributions. While we use Wasserstein Distance due to its performance, the other well-known metrics exhibited comparable results, demonstrating the robustness of our method in terms of the choice of a distance metric and further supporting the effectiveness of our proposed approach.
@@ 5 Conclusion
In this work, we present CASS, a training-free approach for open-vocabulary semantic segmentation that integrates object-level context into CLIP. By distilling object-aware structural information from VFM into CLIP’s attention, CASS effectively enhances intra-object coherence, ensuring that each object is consistently represented as a unified semantic entity. Additionally, we leverage an object presence prior knowledge to adjust text embeddings and refine object class assignments, aligning them with object-specific semantic targets in the image. Consequently, CASS addresses key limitations of training-free OVSS in accurately grouping object components and assigning semantic classes without extra training or additional labeled data, making it a robust solution across diverse datasets and visual contexts.

### v2

@@ 1 Introduction
Open-vocabulary semantic segmentation (OVSS) [33, 57, 10, 56, 60] aims to predict pixel-level labels for arbitrary prompts defined by user input (e.g., the proper noun Space Needle in Fig. 1(a)), advancing beyond previous models scoped to predefined static classes [41, 20, 11, 19]. To achieve this, fully supervised OVSS models [12, 58, 55] enhance generalization to seen classes using labeled training data. Yet, their reliance on seen classes during training may risk overfitting and limit scalability, as labeled data is required for retraining to adapt to new domains [44].
@@ 2 Related Works
Open-vocabulary semantic segmentation (OVSS) aims to segment images corresponding to arbitrary user-given prompts, using vision-language aligned representations (e.g., CLIP [40]). Meanwhile, training-free approaches [4, 5, 30, 60, 24] have risen for their ability to adapt directly to new tasks without extra training. Among these, methods [50, 60, 30, 17] that do not rely on additional datasets have gained attention, as they require no extra adaptation effort, aligning better with the objective of OVSS. Earlier approaches investigated leveraging CLIP features for better localization [60, 31] or for grounding proposed masks [48, 22]. SCLIP [50] uses correlative self-attention to improve inter-patch correlations. NACLIP [17] applies a Gaussian kernel to attention maps for better spatial consistency, and ProxyCLIP [30] integrates patch-wise consistent features from DINO [9]. Although these methods demonstrated strong results, they often struggle to group different parts of the same object due to a lack of explicit dense label knowledge. To this end, we attribute the underlying challenges of training-free OVSS by enriching the object-level contextual understanding.
@@
Object-Level Semantic Segmentation. Semantic segmentation groups object components that have similar semantics into a single mask, traditionally achieved through models trained on extensive human-labeled, pixel-wise annotations [20, 11, 41, 26]. However, despite recent advancements in unsupervised semantic segmentation [18, 43] (USS) that do not rely on dense labels, a core challenge persists: individual object components are not consistently grouped into a single entity. To address this issue, recent USS studies [25, 52, 42] have incorporated object-level representation learning to improve the model’s ability to merge single objects into unified entities. However, these efforts have not been considered in the training-free OVSS task, which faces similar challenges in an unsupervised context. Therefore, we focus on enriching object-level context in training-free OVSS to achieve a more effective grouping of components within each object, enabling clearer and more consistent segmentation across diverse visual domains.
@@ 3.2 Spectral Object-Level Context Distillation
Our model CASS aims to capture such object-level contextual interaction within the image features to associate different parts of the same object. However, the limited ability of CLIP to capture semantic relationships between patches often neglects the object-level context, as illustrated in the left column of Fig. 3 (${A}_{\text{CLIP}}$). To address this limitation, we leverage VFM (e.g., DINO [9]), enabling a deeper object-level contextual understanding. Among VFM features, we utilize VFM attention key component ${K}_{\text{VFM}}\in\mathbb{R}^{h\times N\times D_{h}}$. Specifically, we treat the VFM attention adjacency ${A}_{\text{VFM}}={K}_{\text{VFM}}{K}_{\text{VFM}}^{\top}\in\mathbb{R}^{h\times N\times N}$ as a graph, where $h$ represents the number of heads, and transfer crucial graph patterns from ${A}_{\text{VFM}}$ to ${A}_{\text{CLIP}}$. However, a key question remains: “How do we account for the multi-head attention mechanisms in both VFM and CLIP?" In other words, “how do we match $i$-th VFM attention head ${A}_{\text{VFM}}^{i}$ with $j$-th CLIP attention head ${A}_{\text{CLIP}}^{j}$?" A straightforward approach is to distill the attention heads sequentially (i.e., $i=j$). Nevertheless, recent studies [61, 32, 23] indicate that different attention heads focus on separate parts of the image. Thus, matching the optimal heads between VFM and CLIP is essential for precise attention distillation.
@@ 3.3 Object Presence-Driven Object-Level Context
where $\gamma$ controls the balance between $\hat{\mathcal{S}}$ and object presence prior. This allows the final segmentation map to capture object-level details for precise, class-aware segmentation.
@@ 4.2 Evaluation Results
We evaluate our model against training-free OVSS methods both quantitatively and qualitatively. For fair comparison, we refrain from using the mask refinement (e.g., PAMR [2] or DenseCRF [27]) and reproduce results from [30] using a $224\times 224$ window to match the baselines. Performance is measured primarily with mean IoU (mIoU), complemented by pixel accuracy (pAcc) for a broader assessment.
@@ 4.3 Ablation Study
Distance Metric for Graph Matching. We evaluate different distance metrics $\mathcal{D}_{W}$ used in Sec. 3.2.1, including Wasserstein Distance, Cosine Distance, KL Divergence, and Euclidean Distance. As shown in Fig. 8, the Wasserstein Distance shows the best results as it provides an intuitive measure of the overall shape of sorted eigenvalue distributions. While we use Wasserstein Distance due to its performance, the other well-known metrics exhibited comparable results, demonstrating the robustness of our method in terms of the choice of a distance metric and further supporting the effectiveness of our proposed approach. Acknowledgement. This work was supported in part by the IITP RS-2024-00457882 (AI Research Hub Project), NRF RS-2024-00345806, and NRF RS-2023-00219019.
@@ 5 Conclusion
In this work, we present CASS, a training-free approach for OVSS that integrates object-level context into CLIP. By distilling object-aware structural information from VFM into CLIP’s attention, CASS effectively enhances intra-object coherence, ensuring that each object is consistently represented as a unified semantic entity. Additionally, we leverage object-presence prior knowledge to adjust text embeddings and refine object class assignments, aligning them with object-specific semantic targets in the image. Consequently, CASS effectively groups object components and assigns semantic labels without extra training or additional labeled data, making it a robust solution across diverse datasets and visual contexts.
Supplementary Material
6 Additional Material: Project Page & Presentation Video
We have described our results in an easily accessible manner on our project page, where a brief presentation video is also available. The link to the project page is as follows: https://micv-yonsei.github.io/cass/.
7 Detailed Method
7.1 Energy-based Low-rank Approximation
Input: Adjacency matrix ${A}\in\mathbb{R}^{n\times n}$, energy threshold $\eta$, initial rank $q_{0}$, step size $\Delta q$
Output: Optimal rank $k$, eigenvectors ${U}$, eigenvalues ${\Sigma}$
Set $q_{\text{max}}\leftarrow n$ ;
  // Maximum allowable rank for symmetric matrices
Set $q\leftarrow q_{0}$;
if $\|{A}-{A}^{\top}\|_{F}>\epsilon$ then
       Error: Input matrix is not symmetric;
while $q\leq q_{\text{max}}$ do
       Approximate ${A}$ with rank-$q$ components:
${A}\approx{U}_{q}\Sigma_{q}{U}_{q}^{\top},$
where ${U}_{q}$ contains the top $q$ eigenvectors and $\Sigma_{q}=\text{diag}(\lambda_{1},\lambda_{2},\ldots,\lambda_{q})$;
      Compute total energy: $E_{\text{total}}\leftarrow\sum_{i=1}^{q}\lambda_{i}$
      Compute cumulative energy: $E_{\text{cumulative}}(k)\leftarrow\sum_{i=1}^{k}\lambda_{i}$
      if $\frac{E_{\text{cumulative}}(k)}{E_{\text{total}}}\geq\eta$ for some $k\leq q$ then
             return $k,{U}_{k},\Sigma_{k}$;
      Increment $q\leftarrow q+\Delta q$;
Set $q\leftarrow q_{\text{max}}$ ;
  // Fallback to maximum rank if threshold not met
return $q,{U}_{q},\Sigma_{q}$;
Algorithm 1 Optimal Rank $k$ Selection with Low-Rank Eigendecomposition
We leverage the Energy-based Low-rank Approximation method, as outlined in Algorithm 1, which efficiently captures key object-level contextual features within the VFM graph. The method approximates the $i$-th head attention adjacency graph ${A}_{\text{VFM}}^{i}$ using the rank-$q$ approximation as
${A}_{\text{VFM}}^{i}\approx{U}_{q}\Sigma_{q}{U}_{q}^{\top},$ (10)
where ${U}_{q}$ contains the top $q$ eigenvectors, and $\Sigma_{q}=\text{diag}(\lambda_{1},\lambda_{2},\dots,\lambda_{q})$ consists of the top $q$ eigenvalues. This selective focus on the most significant components highlights critical object relationships while discarding noise. By iteratively increasing the rank $q$, the algorithm efficiently captures the graph’s essential structure with minimal complexity by examining cumulative energy. Our algorithm evaluates whether the retained energy satisfies the threshold $\eta$ as
$\frac{E_{\text{cumulative}}}{E_{\text{total}}}=\frac{\sum_{i=1}^{k}\lambda_{i}}{\sum_{i=1}^{q}\lambda_{i}}\geq\eta,\quad k\leq q.$ (11)
By adaptively identifying the smallest rank $k$ that meets this criterion, the method ensures a compact and meaningful representation. Thus, our low-rank approximated VFM graph is expressed as
$\tilde{A}_{\text{VFM}}^{i}={U}_{k}\Sigma_{k}{U}_{k}^{\top}.$ (12)
Our proposed combination of selective eigendecomposition, energy-based evaluation, and adaptive rank selection enables the algorithm to efficiently preserve object-level contextual features while minimizing computational costs. In the next section, we refine the computed low-rank components (i.e., the eigenvalues and eigenbasis), through dynamic eigenscaling. This approach further enhances their transformation into more object-centric representations, enabling improved precision and robustness in feature modeling.
7.2 Dynamic Eigenscaling
We propose a dynamic eigenscaling function $\phi$ to refine the eigenvalues $\Sigma_{k}$ from Eq. (12). This function amplifies larger eigenvalues while suppressing smaller ones, highlighting dominant components in the graph structure and reducing the influence of noise. The scaled eigenvalues are computed as
$\Sigma_{k}^{\prime}=\phi(\Sigma_{k})=\frac{\Sigma_{k}-\Sigma_{\text{min}}}{\Sigma_{\text{max}}-\Sigma_{\text{min}}}\cdot\text{scaled\_range}+(2-\epsilon)\cdot\Sigma_{\text{min}},$ (13)
[…]
$\ddot{A}_{\text{VFM}}^{i}={U}_{k}\Sigma_{k}^{\prime}{U}_{k}^{\top}:={U}_{k}\phi(\Sigma_{k}){U}_{k}^{\top}.$ (14)
This transformation dynamically adjusts the eigenvalues to retain the graph’s most significant features, enhancing object-level contextual representation while minimizing the impact of less relevant components.
7.3 Hierarchical Grouping of Text Embeddings
from sklearn.metrics.pairwise import cosine_distances
from scipy.cluster.hierarchy import linkage, fcluster
from scipy.spatial.distance import squareform
def hierarchical_clustering(text_embeddings, h_threshold):
# Compute cosine distance matrix
distance_matrix = cosine_distances(text_embeddings)
# Convert distance matrix to condensed form
condensed_matrix = squareform(distance_matrix)
# Perform hierarchical clustering using Ward’s method
Z = linkage(condensed_matrix, method=’ward’)
# Form flat clusters based on the threshold
clusters = fcluster(Z, t=h_threshold, criterion=’distance’) - 1
return clusters
Algorithm 2 Pseudo-code of hierarchical clustering in Python style.
We provide details on grouping class prompts to identify semantically related object class prompts (e.g., motorcycle and bicycle) within the Object-Guided Text Embedding Adjustment module. As described in Algorithm 2, we compute the semantic distance between user-defined text prompts by measuring the cosine distance of their CLIP text embeddings in the CLIP latent space, forming a distance matrix. Next, the distance matrix is converted into a condensed form to facilitate hierarchical clustering using Ward’s method, which minimizes variance within clusters. Finally, clusters are formed by applying a predefined height threshold to the hierarchical tree, ensuring that semantically similar class prompts are grouped together. This enables the identification of class prompt-level relationships that are leveraged to refine the text embeddings towards object-specific semantics.
8 Additional Experiments
8.1 Additional Evaluation Results
In this section, we provide additional evaluation results including qualitative evaluation (8.1.1), real-world open-vocabulary semantic segmentation result (10.2), and scale-up version of CASS (8.1.2).
8.1.1 Additional Qualitative Evaluation
We present additional qualitative evaluation results: Fig. 11 illustrates results on PASCAL VOC (V20 & V21) [16], Fig. 12 on COCO (C-Stf & C-Obj) [8], Fig. 13 on PASCAL Context (PC59 & PC60) [38], and Fig. 14 on both ADE20K (ADE) [59] and Cityscapes (City) [14]. Note that all the results in this subsection are presented without mask refinement steps, such as PAMR [1] or DenseCRF [27], to ensure a fair comparison and to evaluate the capability of the model itself.
8.1.2 Scale-up Version
Table 5: Performance of scale-up version of CASS.
Model Feature Encoder [CLS] Encoder V21 PC60 C-Obj V20 PC59 ADE Avg
CaR [48] CVPR’24 ViT-B/16 ViT-B/16 48.6 13.6 15.4 73.7 18.4 5.4 29.2
CASS ViT-B/16 ViT-B/16 65.8 36.7 37.8 87.8 40.2 20.4 48.1
CaR [48] CVPR’24 ViT-B/16 ViT-L/14 63.7 29.4 35.1 91.4 38.4 16.9 45.8
CASS ViT-B/16 ViT-L/14 66.3 37.0 38.3 89.3 40.7 20.7 48.7
The zero-shot object classification performance of CLIP [40] improves as the capacity of CLIP increases. Unlike the main experiment, where all methods were standardized using CLIP ViT-B/16 for a fair comparison, here we compute the object presence prior using CLIP ViT-L/14 (i.e., the scaled-up version), as reported in Table 5. Since CaR [48] also classifies the proposed mask using CLIP’s global image embedding for classification, we compared our method under the same conditions as CaR. The encoder used for patch-wise dense representation (i.e., $F_{\text{CLIP}}$ in our main manuscript) is referred to as the Feature Encoder, while the encoder for encoding CLIP visual embedding vector (i.e., $v_{\text{CLIP}}$) is referred to as the [CLS] Encoder. The CLIP visual embedding vector serves as an object presence prior in our approach and as a mask proposal classifier in CaR. For this experiment, we use datasets that are reported in CaR and evaluate with the mIoU metric. Note that we exclude a mask refinement step (e.g., PAMR [1] or DenseCRF [27]) for a fair comparison.
When using CLIP ViT-L/14 as the [CLS] encoder in our CASS, we observe a noticeable performance improvement compared to the ViT-B/16 configuration. Furthermore, our approach consistently outperforms CaR across all benchmarks under both encoder settings. Notably, when comparing models fairly by using the same [CLS] encoder configuration as CaR (ViT-B/16), CaR experiences a significant drop in performance, particularly on datasets such as PC60 and ADE. In contrast, our method demonstrates robust performance in both settings, maintaining high accuracy and adaptability regardless of the encoder used. This highlights the stability and effectiveness of our method, especially in handling variations in encoder configurations, demonstrating its robustness compared to CaR.
8.2 Additional Ablation Study
For ablation studies, we use representative datasets for segmentation tasks: PASCAL VOC (V21), PASCAL Context (PC59), and COCO-Stuff (C-Stf), following [50].
8.2.1 Ablation: VFM Backbones
Table 6: Ablation results with different VFM backbones. For the CLIP visual encoder, ViT-B/16 is used.
[…]
mIoU pAcc mIoU pAcc mIoU pAcc
[…]
ProxyCLIP [30] ECCV’24 59.1 86.6 38.8 63.4 26.2 43.4 41.4 64.5
CASS 65.8 90.1 40.2 65.0 26.7 43.6 44.2 66.2
[…]
ProxyCLIP [30] ECCV’24 56.6 85.8 37.4 62.0 25.1 42.2 39.7 63.3
CASS 64.3 89.1 38.9 63.8 26.1 43.1 43.1 65.3
[…]
ProxyCLIP [30] ECCV’24 57.1 85.2 37.3 61.4 25.3 42.3 39.9 63.0
CASS 63.0 88.6 38.1 62.2 25.3 42.0 42.1 64.3
[…]
Table 6 presents the results with various VFM backbones [9, 39], including DINOv1 ViT-B/8, DINOv1 ViT-B/16, and DINOv2 ViT-B/14. Among these, DINOv1 ViT-B/8 serves as the backbone for our main results. In this experiment, we use ViT-B/16 as the CLIP visual encoder. To ensure a consistent and fair comparison, we include LaVG [22] and ProxyCLIP [30] as baselines, as both utilize DINO as their backbone model. Since LaVG does not report performance on DINOv1 ViT-B/16 and DINOv2 ViT-B/14, we report reproduced results. Notably, LaVG employs spectral techniques, particularly graph partitioning, making it directly comparable to our approach, which also builds on spectral methods. The results show that our method consistently outperforms the baselines across all backbones, demonstrating its robustness and adaptability to different VFMs. In particular, compared to LaVG, which uses graph partitioning for spectral techniques, our method more effectively leverages spectral methods by distilling low-rank components into CLIP via spectral-based graph matching, offering a more advanced solution for training-free OVSS.
8.2.2 Ablation: Feature Type
Table 7: Comparison of features used for constructing attention graphs.
[…]
mIoU pAcc mIoU pAcc mIoU pAcc
[…]
$FF$ $QK$ 44.6 77.3 30.2 54.0 19.9 37.1 31.6 56.1
$FF$ $KK$ 60.8 87.6 37.9 62.1 25.2 42.0 41.3 63.9
$QQ$ $QQ$ 57.1 84.9 36.5 61.2 23.9 41.5 39.2 62.5
$QQ$ $QK$ 50.9 53.1 33.7 43.6 21.9 30.2 35.5 42.3
$QQ$ $KK$ 64.9 89.3 39.7 64.5 26.4 43.5 43.7 65.8
$QK$ $QQ$ 33.7 58.1 25.2 45.7 17.3 32.4 25.4 45.4
$QK$ $QK$ 30.7 53.1 23.2 43.6 15.5 30.2 23.1 42.3
$QK$ $KK$ 29.0 49.1 22.1 41.7 15.0 28.6 22.0 39.8
$KK$ $QQ$ 64.9 89.3 40.0 64.6 26.6 43.7 44.0 65.8
$KK$ $QK$ 60.3 87.6 38.5 63.4 25.4 43.0 41.4 64.7
$KK$ $KK$ 65.8 90.1 40.2 65.0 26.7 43.6 44.2 66.2
[…]
Table 7 presents the performance impact of various features employed in constructing $A_{\psi}$ during VFM graph distillation. The VFM column represents the features utilized in the construction of $A_{\text{VFM}}$, while the CLIP column represents the features used for $A_{\text{CLIP}}$. In the table, $F$ refers to visual features derived from VFM, and $Q$ and $K$ represent the attention query and key components from their respective models. Each row represents a specific combination of features from VFM and CLIP used to construct the attention graph (e.g., $QK$ in the VFM column indicates that $A_{\text{VFM}}$ is constructed as $A_{\text{VFM}}=QK$).
When using visual features $F$ from VFM, the results show moderate performance overall; however, constructing $A_{\text{VFM}}$ using attention components such as $Q$ or $K$ instead of $F$ resulted in better performance. This indicates that attention components more effectively capture semantic relationships between patches, demonstrating the efficiency of using attention to extract object-level context, which aligns with the findings of prior study [25]. In terms of attention components, $QQ$ performs better than $QK$ across most datasets but is outperformed by $KK$. The $QK$ configuration generally shows weaker performance compared to both $QQ$ and $KK$, highlighting the importance of modifying the last layer of self-attention in CLIP instead of using vanilla attention settings. Overall, the best performance is achieved when $KK$ is used for the attention graph consistently in both VFM and CLIP, as demonstrated by the highest average mIoU of 44.2 and average pAcc of 66.2. This suggests that the $KK$ provides the most robust and object-centric representation for constructing attention graphs.
8.2.3 Ablation: Graph Matching Strategy
Table 8: Ablation results on different graph matching strategies.
[…]
mIoU pAcc mIoU pAcc mIoU pAcc
[…]
Similar 65.5 90.0 40.2 65.1 26.6 43.6 44.1 66.2
Complementary 65.8 90.1 40.2 65.0 26.7 43.6 44.2 66.2
[…]
In Table 8, we compare different graph matching strategies. “Sequence" refers to matching graphs head-to-head in a sequential manner without applying additional structural alignment techniques. “Similar" involves matching graphs that exhibit similar structures between VFM and CLIP. Lastly, “Complementary" uses a strategy that matches graphs with contrasting but complementary structures between VFM and CLIP to leverage diverse information. From the results, we observe that the “Sequential" method delivers competitive performance, achieving an average mIoU of 43.5 and pAcc of 65.7. The rows for “Similar” and “Complementary” consider the attention graph structures from each head, matching either similar or dissimilar ones, respectively. The “Similar" strategy improves the results, with an average mIoU of 44.1 and pAcc of 66.2. However, the proposed “Complementary" method outperforms both matching methods, achieving the highest average mIoU of 44.2 and pAcc of 66.2. Overall, while sequential approaches are straightforward and offer reasonable performance, these results highlight the advantages of considering in matching attention heads.
8.2.4 Ablation: Number of Eigenvalues for Low-Rank Components in VFM Graph
Table 9: Ablation results on rank selection methods for low-rank components.
[…]
mIoU pAcc mIoU pAcc mIoU pAcc
[…]
$k=4$ 64.5 89.8 39.8 64.6 26.4 43.1 43.6 65.8
Eigengap 64.9 89.6 39.8 64.6 26.3 43.3 43.7 65.8
Energy-based 65.8 90.1 40.2 65.0 26.7 43.6 44.2 66.2
[…]
In this experiment, we compare methods for selecting the rank (number of eigenvalues) for low-rank components. Table 9 presents the results of applying several well-known methods alongside our proposed energy-based approach. Rows marked with ✗ represent the use of a fixed rank $k$ across the entire images in datasets, whereas rows marked with ✓ indicate searching for the optimal rank on a per-image basis. The eigengap is computed as the index $i^{*}$ where the difference between consecutive eigenvalues in the spectrum is maximized, typically defined as $i^{*}=\underset{i\in\{1,2,\dots,n-1\}}{\arg\max}\,(\lambda_{i}-\lambda_{i+1})$, where $\lambda_{i}$ and $\lambda_{i+1}$ represent the $i$-th and $(i+1)$-th largest eigenvalues, respectively. This metric is commonly used to identify the point where the eigenvalue spectrum exhibits a significant drop, which helps in determining the optimal rank for low-rank approximations or clustering applications.
Fixed-rank methods maintain consistent ranks across all images but fail to adapt to varying data complexity within individual images. As a result, their performance is lower compared to adaptive approaches. The eigengap method, which adaptively selects the rank by identifying a significant drop in the eigenvalue spectrum, improves performance compared to fixed-rank strategies. However, the eigengap approach is limited by its sensitivity to small variations in the eigenvalue distribution, which can lead to inconsistent rank selections in certain cases. Our energy-based method further enhances performance by selecting ranks based on the cumulative energy of the eigenvalues, ensuring a more stable and data-driven approach to rank determination. This method achieves the best overall results and consistently outperforms all other methods across individual datasets. These results highlight the importance of adaptive rank selection strategies in capturing the intrinsic object-level structure of graphs.
9 Implementation Details
9.1 Prompt Templates
Table 10: Examples of the prompt templates used for text embedding generation.
Examples of Prompt Templates
[…]
9.2 Hyperparameters
Table 11: Hyperparameters used in CASS.
V21 PC60 C-Obj V20 PC59 C-Stf City ADE
$\alpha$ $\gamma$ $\alpha$ $\gamma$ $\alpha$ $\gamma$ $\alpha$ $\gamma$ $\alpha$ $\gamma$ $\alpha$ $\gamma$ $\alpha$ $\gamma$ $\alpha$ $\gamma$
0.03 0.10 0.04 0.25 0.01 0.25 0.02 0.40 0.04 0.25 0.02 0.20 0.03 0.10 0.05 0.30
Here, we provide a detailed description of the hyperparameters listed in Table 11. The parameter $\alpha$ controls the balance between the original text embedding and the object-specific vector, while $\gamma$ adjusts the trade-off between the patch-text similarity $\hat{\mathcal{S}}$ and the object presence prior.
10 Application
10.1 Image Inpainting and Object Removal
Figure 9: Visualization of image inpainting and object removal using our predicted mask. For image inpainting, we use “red sports car with red wheels" as an input prompt. Note that the mask refinement step is excluded when segmenting the object mask.
While our CASS can naturally be used for semantic segmentation directly based on user-provided prompts in real-world scenarios (see Sec. 10.2 for diverse examples), its ability to be object-level context-aware allows it to group object components effectively. As a result, CASS produces clean object masks that can be effectively utilized in downstream tasks (e.g., image inpainting and object removal). Fig. 9 visualizes the application of our CASS in various downstream tasks. When provided with the object prompt Toyota car, our model generates a precise object mask, enabling tasks such as image inpainting and object removal. For image inpainting, we utilize Stable Diffusion XL (SDXL), while for object removal, we employ LaMa [49]. The baseline model used for comparison is NACLIP [17]. Comparing the results to the baseline, it is evident that the baseline model struggles to correctly mask all the object components, such as the rear wheels and headlights. For example, the baseline fails to mask the rear wheel of the car, which limits the inpainting model from accurately rendering the red wheel in the edited image. Similarly, incomplete masking of headlights results in unedited regions (e.g., headlight) during the inpainting process. This limitation also affects object removal tasks. The baseline’s failure to mask the entire object (i.e., Toyota car) limits object-level removal and instead focuses on specific internal components, such as side mirrors or door handles. In contrast, our method successfully captures all relevant object components, producing a complete and accurate object mask that enables the removal of the object as a whole. These results demonstrate the importance of generating object-level context-aware masks, as they enable downstream models to perform tasks such as image inpainting and object removal with significantly improved precision and quality.
10.2 Open-Vocabulary Semantic Segmentation in the Wild
Fig. 15 illustrates the application of CASS on real-world images, demonstrating its ability to adapt to diverse and uncontrolled settings. This showcases the method’s robustness in handling the complexities and variations typically encountered in real-world scenarios, without requiring additional fine-tuning or preprocessing.
11 Discussion
11.1 Limitations
Figure 10: Visualization of limitation of CASS.
Our CASS introduces object-level contextual knowledge into training-free OVSS, enabling the effective grouping of object components into coherent semantic entities. By combining Energy-based Low-rank Approximation with spectral-based graph matching, our approach narrows the gap between pixel-level predictions and object-level understanding, contributing a significant advancement towards the primary objectives of OVSS.
However, CASS has limitations, particularly in computational efficiency. While Energy-based Low-rank Approximation reduces costs, the eigendecomposition step remains computationally demanding, especially for high-resolution images. Additionally, the Hungarian matching algorithm, which runs on the CPU, introduces further latency due to its lack of GPU support. These inefficiencies hinder real-time applicability. For instance, CASS processes at 5.6 FPS on an RTX A6000 GPU, which is insufficient for real-time segmentation. Optimizing computational efficiency is essential to expanding its practical use in time-sensitive tasks.
Moreover, as shown in Fig. 10, our emphasis on object-level context leads to stronger performance on large objects (e.g., building: 20.0 [30] to 34.7 mIoU) but can underperform on smaller objects, ultimately diminishing the overall mIoU improvement due to the averaging across all classes. As a result, CASS yields substantial improvements on datasets with predominantly large and distinct object classes (e.g., VOC), while showing relatively modest gains on datasets rich in small object categories (e.g., ADE, COCO). Nonetheless, CASS still secures a 7.25% relative improvement over the existing SoTA [30] on average across 8 datasets, underscoring its meaningful contribution to advancing training-free OVSS.
11.2 Future Works
Our proposed CASS is designed to perform semantic segmentation on individual images based on user-provided arbitrary prompts, achieving competitive performance in grouping object components into coherent entities. While effective for static scenarios, the current approach does not account for temporal consistency, limiting its applicability to video sequences and dynamic environments. To extend its use to real-time or sequential analysis tasks, future work will focus on incorporating temporal information to ensure coherence across frames. This enhancement will enable the method to handle video streams more effectively, supporting applications such as object tracking and dynamic scene understanding.
As mentioned in the limitations, the high computational time remains a challenge that must be addressed as part of our future work. Specifically, optimizing the eigendecomposition process and the Hungarian matching algorithm used for graph matching is crucial. These improvements will play a key role in enhancing the efficiency of our method and making real-time inference feasible for practical applications.
Figure 11: Additional qualitative comparison between recent state-of-the-art methods SCLIP [50], NACLIP [17], ProxyCLIP [30] and Our CASS using PASCAL VOC dataset [16].
Figure 12: Additional qualitative comparison between recent state-of-the-art methods SCLIP [50], NACLIP [17], ProxyCLIP [30] and Our CASS using COCO dataset [8].
Figure 13: Additional quantitative comparison between recent state-of-the-art methods SCLIP [50], NACLIP [17], ProxyCLIP [30] and Our CASS using PASCAL Context [38].
Figure 14: Additional qualitative comparison between recent state-of-the-art methods SCLIP [50], NACLIP [17], ProxyCLIP [30] and Our CASS using ADE20K [59] (top) and Citscapes [14] (bottom).
Figure 15: Real-world open-vocabulary semantic segmentation results of our model after applying mask refinement [1].